# No-Regret Strategy Solving in Imperfect-Information Games via Pre-Trained Embedding

### v1

@@ 3.3 Convergence Analysis
Second, even in the advisor space, proving that the embedded cumulative regret of a single advisor stably converges to zero remains challenging. This is because each infoset is simultaneously shaped by all advisors — a dynamic where their update processes are mutually interdependent, with adjustments to one advisor rippling through and altering the convergence trajectory of others.
Fortunately, we can show that when an advisor acts in isolation, Embedding CFR ensures its regret decreases. Embedding CFR can be conceptualized as follows: in iteration $T$, when player $i$ encounters an infoset $I_{q}\in J$ ($J$ is an info-block relevant to player $i$) during the game, a random selection is made to decide with probability $\phi_{p,q}$ to act according to the strategy $\sigma^{T}_{i}(e_{p})$. Consider an approximate sampling scenario for Embedding CFR: in each iteration, we strictly need to consider all infosets in $J$. The key difference is that for each infoset, a selection must be made among $e_{1},\dots,e_{m}$ — choosing to interact and update according to a specific advisor’s strategy. At a particular iteration $T$, it happens that player $i$ selects $e_{p}$ for every infoset $I_{q}$. This process is illustrated by the green and red arrows in Figure 2, where all infosets have selected advisor $e_{2}$. This process approximates the impact of an advisor acting in isolation on its regret within Embedding CFR.
@@
Let $S^{T}_{p}=\sum_{a\in A(J)}\left(R^{T}(e_{p},a)_{+}\right)^{2}$ and $\Delta_{J}=\max_{\begin{subarray}{c}\sigma\in\Sigma\\ I\in J\\ a,a^{\prime}\in A(J)\end{subarray}}\bigl|v_{\mathcal{P}(I)}^{\sigma_{|I\to a}}(I)-v_{\mathcal{P}(I)}^{\sigma_{|I\to a^{\prime}}}(I)\bigr|$. In the aforementioned scenario, the following holds: If $S^{T}_{p}\leq\frac{n|A(J)|\cdot\Delta_{J}^{2}\cdot\|\mathbf{\Phi}_{p,\cdot}\|_{2}^{2}}{T}$, then $S^{T+1}_{p}\leq\frac{n|A(J)|\cdot\Delta_{J}^{2}\cdot\|\mathbf{\Phi}_{p,\cdot}\|_{2}^{2}}{T+1}$. Otherwise, $S^{T+1}_{p}\leq\frac{T}{T+1}S^{T}_{p}$.
@@ 4.1 Partitioning of Info-blocks by Chance Actions
On the other hand, algorithmically, Embedding CFR operates per info-block, requiring a embedding matrix $\mathbf{\Phi}$ for each — a labor-intensive process. If partitions allowed matrices to be constructed for a subset of info-blocks, with others extendable from these, workload would drop markedly.
@@
all the three histories (and their containing infosets $I,I^{\prime}$) belong to the same info-block $J$ by our partitioning rule, as they share the identical non-chance trace: $\mathcal{H}_{-c}(h)=\mathcal{H}_{-c}(h^{\prime})=\mathcal{H}_{-c}(h^{\prime\prime})=\emptyset\cdot\texttt{rRc}\cdot\emptyset\cdot\texttt{Cr},$ where $\mathcal{H}_{-c}(\cdot)$ denotes the operator that retains only non-chance actions (replacing chance segments with $\emptyset$).
Furthermore, this partitioning rule directly addresses the challenge of constructing embedding matrices for each info-block. For example, consider histories $\dot{h},\dot{h}^{\prime},\dot{h}^{\prime\prime}$ with the same chance traces as $h,h^{\prime},h^{\prime\prime}$ but a different non-chance trace ($\emptyset\cdot\texttt{rC}\cdot\emptyset\cdot\texttt{C}$) — the infosets to which these histories belong come from another info-block $\dot{J}$:
@@
Notably, $J$ and $\dot{J}$ are in one-to-one correspondence: their infosets (e.g., $\leftrightarrow\dot{I}, II$$^{\prime}\leftrightarrow\dot{I}^{\prime}$) are counterparts linked by identical chance traces. Since history indistinguishability within poker infosets stems solely from such chance actions (not non-chance actions), we construct embedding matrix $\mathbf{\Phi}$ based on chance actions (i.e., card dealing) — details in Section 4.2. Given that $J$ and $\dot{J}$ share this chance-based structure via their corresponding infosets, $\mathbf{\Phi}$ built for $J$ can be directly reused for $\dot{J}$, substantially reducing construction workload.
@@ 4.2 Embedding Matrices: Construction and Deployment in Poker Games
Functionally, HandEbdNet comprises three core components: (1) a hand feature encoder (HFE) — a composite module of multiple networks — that maps $\mathbf{x}_{q}$ to an m-dimensional intermediate feature vector; (2) a softmax layer transforming this vector into an m-dimensional probability distribution, viewed as $\mathbf{\Phi}_{\cdot,q}$; and (3) a final fully connected layer (FC) that transforms this distribution, with supervision from $\mathbf{y}_{q}$ forming the learning objective.
The core insight is using $\mathbf{y}_{q}$ to induce discriminative $\mathbf{\Phi}_{\cdot,q}$: hands with similar strength characteristics (reflected in $\mathbf{y}_{q}$) yield proximate $\mathbf{\Phi}_{\cdot,q}$, capturing meaningful structural relationships between infosets in the advisor space. In poker scenarios, traversing all possible hands (rather than storing them) is feasible, and no entirely new hands will be encountered during embedding coordinate generation. Thus, HandEbdNet is not prone to overfitting in both training and deployment. Details on constructing $\mathbf{x}_{q}$, $\mathbf{y}_{q}$, and HandEbdNet’s full architecture are provided in Appendix C.
@@ 5 Experiment
We conducted experiments in the Numeral211 Hold’em environment, a simplified variant of Texas Hold’em poker proposed by Fu et al. (2024). This game uses a 40-card deck and includes 3 betting rounds, retaining sufficient complexity—particularly in hand diversity—while reducing the overall game scale. This makes it an ideal testbed for researching information set abstraction. Detailed rules are provided in Appendix A.3.
We primarily compare strategies generated by (extended) information set abstraction methods—including the classical EHS, PaEmd (an algorithm successfully applied in DeepStack and Libratus), and the novel KrwEmd proposed by (Fu et al. 2024)—when paired with vanilla CFR for solving, alongside our Embedding CFR, which simultaneously handles information set abstraction and strategy solving.
To ensure a fair comparison, all algorithms are maintained at the logical spatial resources (in terms of the number of abstracted information sets or the size of the advisor space). For the Numeral211 Hold’em environment, the number of hand combinations across its three betting rounds is $\binom{40}{2}$, $\binom{40}{2}\times\binom{38}{1}$, and $\binom{40}{2}\times\binom{38}{1}\times\binom{37}{1}$ respectively. The number of structurally equivalent hand strength classes (via lossless isomorphism (Gilpin and Sandholm 2007b)) across the three rounds is 100, 2250, and 3957, as determined by Waugh (2013). Regarding the logical spatial resources used by abstraction algorithms and Embedding CFR, they are restricted to 225 and 396 in betting rounds 2 and 3 respectively. Notably, all algorithms do not apply abstraction or embedding in the first round; instead, they rely on lossless isomorphism.
@@
We compare the convergence of exploitability among these algorithms, and the experimental results are presented in Figure 5. For the experiment, we sampled a number of hands equivalent to all possible hand combinations in a game, which is $\binom{40}{2}\times\binom{38}{1}\times\binom{37}{1}$ (approximately $10^{6}$) as one iteration, and a total of 1024 iterations are conducted to compare the exploitability in various scenarios. The unit of exploitability is milli blind per game (mb/g). For each baseline comparison algorithm, we provide three parameter configurations. The figure shows the optimal experimental curve of the algorithm. More extensive experiments and details of the experimental equipment are provided in Appendix D.
@@ 6 Conclusion
In conclusion, this paper presents the Embedding CFR algorithm, which innovatively introduces the embedding concept to address the pre-training process of information set abstraction in imperfect information games. This approach marks a significant breakthrough in the field, as it comprehensively outperforms clustering-based information set abstraction algorithms.
@@ A.1 Heads-Up Texas Hold’em
    - Employs a standard 52-card deck (excluding Jokers), with each player privately dealt two hole cards preflop.
    - Five community cards are revealed sequentially: three cards (the flop) are dealt first, followed by the turn (fourth card) and river (fifth card), forming the strongest possible 5-card hand in combination with the player’s hole cards.
@@ A.2 Kuhn Poker
    Hole Card: Each player is dealt one private card face down.
@@ A.3 Numeral211 Hold’em
Numeral211 Hold’em is a poker variant more complex than Kuhn Poker but significantly simpler than HULHE. With diverse hand possibilities, it serves as an ideal testbed for studying hand abstraction tasks (Fu et al. 2024).The game proceeds as follows:
@@
    - Each player receives one private hole card preflop. Two community cards are revealed sequentially, with one card in flop and one card in turn, respectively. The strongest 3-card hand is formed using the hole card and community cards.
@@
Appendix B Theoretical Proofs
B.1 Proof of Proposition 1
Lemma 1 ((Lanctot et al. 2009), Lemma 7).
@@ 2. Large $S^{T}$: Uncontrolled Polynomial Decay (Fractional Scaling)
Appendix C Detailed Implementation of HandEbdNet in Numeral211 Hold’em
C.1 The Construction of the Hand Tensor $\mathbf{x}_{q}$
@@
C.2 The Construction of the Hand Strength Tensor $\mathbf{y}_{q}$
Let’s formalize the construction process of the hand strength tensor $\mathbf{y}$ of a hand $x$. We largely draw on the ideas by Fu et al. (2024) to record the strengths of $x$ and its predecessors in their respective rounds.

### v2

@@ 3.3 Convergence Analysis
Second, even in the advisor space, proving that the embedded cumulative regret of a single advisor stably converges to zero remains challenging. This is because each infoset is simultaneously shaped by all advisors–a dynamic where their update processes are mutually interdependent, with adjustments to one advisor rippling through and altering the convergence trajectory of others.
Fortunately, we can show that when an advisor acts in isolation, Embedding CFR ensures its regret decreases. Embedding CFR can be conceptualized as follows: in iteration $T$, when player $i$ encounters an infoset $I_{q}\in J$ ($J$ is an info-block relevant to player $i$) during the game, a random selection is made to decide with probability $\phi_{p,q}$ to act according to the strategy $\sigma^{T}_{i}(e_{p})$. Consider an approximate sampling scenario in Embedding CFR, each iteration strictly needs to consider all infosets in $J$; and the key difference is that for each infoset, a selection must be made among $e_{1},\dots,e_{m}$–choosing to interact and update according to a specific advisor’s strategy. At a particular iteration $T$, it happens that player $i$ selects $e_{p}$ for every infoset $I_{q}$ (as in Figure 2’s green/red arrows, where all infosets choose $e_{2}$); this approximates an advisor’s isolated impact on its regret.
@@
Let $S^{T}_{p}=\sum_{a\in A(J)}\left(R^{T}(e_{p},a)_{+}\right)^{2}$ and $\Delta_{J}=\max_{\begin{subarray}{c}\sigma\in\Sigma\\ I\in J\\ a,a^{\prime}\in A(J)\end{subarray}}\bigl|v_{\mathcal{P}(I)}^{\sigma_{|I\to a}}(I)-v_{\mathcal{P}(I)}^{\sigma_{|I\to a^{\prime}}}(I)\bigr|$. In the aforementioned scenario, the following holds: If $S^{T}_{p}\leq\frac{n|A(J)|\cdot\Delta_{J}^{2}\cdot\|\mathbf{\Phi}_{p,\cdot}\|_{2}^{2}}{T}$, then $S^{T+1}_{p}\leq\frac{n|A(J)|\cdot\Delta_{J}^{2}\cdot\|\mathbf{\Phi}_{p,\cdot}\|_{2}^{2}}{T+1}$. Otherwise, $S^{T+1}_{p}\leq\frac{T}{T+1}S^{T}_{p}$ (see Appendix C for the proof).
@@ 4.1 Partitioning of Info-blocks by Chance Actions
On the other hand, algorithmically, Embedding CFR operates per info-block, requiring an embedding matrix $\mathbf{\Phi}$ for each–a labor-intensive process. If partitions allowed matrices to be constructed for a subset of info-blocks, with others extendable from these, workload would drop markedly.
@@
all three histories (and their containing infosets $I,I^{\prime}$) belong to the same info-block $J$ by our partitioning rule, as they share the identical non-chance trace: $\mathcal{H}_{-c}(h)=\mathcal{H}_{-c}(h^{\prime})=\mathcal{H}_{-c}(h^{\prime\prime})=\emptyset\cdot\texttt{rRc}\cdot\emptyset\cdot\texttt{Cr},$ where $\mathcal{H}_{-c}(\cdot)$ denotes the operator that retains only non-chance actions (replacing chance segments with $\emptyset$).
Furthermore, this partitioning rule directly addresses the challenge of constructing embedding matrices for each info-block. For example, consider histories $\dot{h},\dot{h}^{\prime},\dot{h}^{\prime\prime}$ with the same chance traces as $h,h^{\prime},h^{\prime\prime}$ but a different non-chance trace ($\emptyset\cdot\texttt{rC}\cdot\emptyset\cdot\texttt{C}$)–the infosets to which these histories belong come from another info-block $\dot{J}$:
@@
Notably, $J$ and $\dot{J}$ are in one-to-one correspondence: their infosets (e.g., $\leftrightarrow\dot{I}, II$$^{\prime}\leftrightarrow\dot{I}^{\prime}$) are counterparts linked by identical chance traces. Since history indistinguishability within poker infosets stems solely from such chance actions (not non-chance actions), we construct embedding matrix $\mathbf{\Phi}$ based on chance actions (i.e., card dealing)–details in Section 4.2. Given that $J$ and $\dot{J}$ share this chance-based structure via their corresponding infosets, $\mathbf{\Phi}$ built for $J$ can be directly reused for $\dot{J}$, substantially reducing construction workload.
@@ 4.2 Embedding Matrices: Construction and Deployment in Poker Games
Functionally, HandEbdNet comprises three core components: (1) a hand feature encoder (HFE)–a composite module of multiple networks–that maps $\mathbf{x}_{q}$ to an m-dimensional intermediate feature vector; (2) a softmax layer transforming this vector into an m-dimensional probability distribution, viewed as $\mathbf{\Phi}_{\cdot,q}$; and (3) a final fully connected layer (FC) that transforms this distribution, with supervision from $\mathbf{y}_{q}$ forming the learning objective.
The core insight is using $\mathbf{y}_{q}$ to induce discriminative $\mathbf{\Phi}_{\cdot,q}$: hands with similar strength characteristics (reflected in $\mathbf{y}_{q}$) yield proximate $\mathbf{\Phi}_{\cdot,q}$, capturing meaningful structural relationships between infosets in the advisor space. In poker scenarios, traversing all possible hands (rather than storing them) is feasible, and no entirely new hands will be encountered during embedding coordinate generation. Thus, HandEbdNet is not prone to overfitting in both training and deployment. Details on constructing $\mathbf{x}_{q}$, $\mathbf{y}_{q}$, and HandEbdNet’s full architecture are provided in Appendix D .
@@ 5 Experiment
We conducted experiments in the Numeral211 Hold’em environment, a simplified variant of Texas Hold’em poker proposed by fu2024signal. This game uses a 40-card deck and includes 3 betting rounds, retaining sufficient complexity—particularly in hand diversity—while reducing the overall game scale. This makes it an ideal testbed for researching information set abstraction. Detailed rules are provided in Appendix A.3 .
We primarily compare strategies generated by (extended) information set abstraction methods—including the classical EHS, PaEmd (an algorithm successfully applied in DeepStack and Libratus), and the novel KrwEmd proposed by fu2024signal—when paired with Vanilla CFR for solving, alongside our Embedding CFR, which simultaneously handles information set abstraction and strategy solving.
To ensure a fair comparison, all algorithms are maintained at the logical spatial resources (in terms of the number of abstracted information sets or the size of the advisor space). For the Numeral211 Hold’em environment, the number of hand combinations across its three betting rounds is $\binom{40}{2}$, $\binom{40}{2}\times\binom{38}{1}$, and $\binom{40}{2}\times\binom{38}{1}\times\binom{37}{1}$ respectively. The number of structurally equivalent hand strength classes (via lossless isomorphism (gilpin2007lossless)) across the three rounds is 100, 2250, and 3957, as determined by waugh2013fast. Regarding the logical spatial resources used by abstraction algorithms and Embedding CFR, they are restricted to 225 and 396 in betting rounds 2 and 3 respectively. Notably, all algorithms do not apply abstraction or embedding in the first round; instead, they rely on lossless isomorphism.
@@
We compare the convergence of exploitability among these algorithms, and the experimental results are presented in Figure 5. For the experiment, we sampled a number of hands equivalent to all possible hand combinations in a game, which is $\binom{40}{2}\times\binom{38}{1}\times\binom{37}{1}$ (approximately $10^{6}$) as one iteration, and a total of 1024 iterations are conducted to compare the exploitability in various scenarios. The unit of exploitability is milli blind per game (mb/g). For each baseline comparison algorithm, we provide three parameter configurations. The figure shows the optimal experimental curve of the algorithm. More extensive experiments and details of the experimental equipment are provided in Appendix E .
@@ 6 Conclusion
In conclusion, this paper presents the Embedding CFR algorithm, which innovatively introduces the embedding concept to address the pre-training process of information set abstraction in imperfect information games. This approach marks a significant breakthrough in the field, as it comprehensively outperforms clustering-based methods in our experimental settings.
Acknowledgments
This work was supported by the National Science and Technology Major Project (Grant No. 2022ZD0116403) and the China Postdoctoral Science Foundation (Grant No. 2024M763533).
@@ A.1 Heads-Up Texas Hold’em
    - Employs a standard 52-card deck (excluding Jokers), with each player privately dealt two private cards preflop.
    - Five community cards are revealed sequentially: three cards (the flop) are dealt first, followed by the turn (fourth card) and river (fifth card), forming the strongest possible 5-card hand in combination with the player’s private cards.
@@ A.2 Kuhn Poker
    Private Card: Each player is dealt one private card face down.
@@ A.3 Numeral211 Hold’em
Numeral211 Hold’em is a poker variant more complex than Kuhn Poker but significantly simpler than HULHE. With diverse hand possibilities, it serves as an ideal testbed for studying hand abstraction tasks (fu2024signal).The game proceeds as follows:
@@
    - Each player receives one private card preflop. Two community cards are revealed sequentially, with one card in flop and one card in turn, respectively. The strongest 3-card hand is formed using the private card and community cards.
@@
Appendix B Pseudocode for Poker-Specific Embedding CFR
Algorithm 1 presents the pseudocode for Poker-Specific Embedding CFR (e.g., tailored for Texas Hold’em), which leverages the unique properties of poker-like games as described in Section 4.1. Specifically, the progression of such games can be partitioned into a sequence of public tree nodes, which depict the game process identifiable to an observer. These nodes correspond to key game events, such as the chance player dealing community cards or other players taking actions like calling, betting, or folding. We denote the set of these public tree nodes as $V$.
All stochastic factors in the game can be determined prior to traversing the entire public tree—these factors refer to the private cards dealt to players and the community cards to be revealed by the final showdown stage. Traversing the public tree with these pre-determined cards is referred to as chance sampling.
Each hand of cards, when combined with a given public tree node $v$, uniquely defines an information set. It is straightforward to confirm that there exists a one-to-one correspondence between each public tree node $v$ and an info-block $J$. We use $J(v)$ to represent the info-block determined by node $v$.
Algorithm 1 Embedding CFR algorithm for two-player hold’em games
[…]
4:for $t\leftarrow 1$ to $n$ do
5:  Deal $l$ batches of final private card tensors $\mathbf{h}_{1},\mathbf{h}_{2}$ (for players 1 and player 2, respectively) and a final community card tensor $\mathbf{b}$
6:  Traverse($v^{0},\mathbf{h}_{1},\mathbf{h}_{2},\mathbf{b},\mathbf{1}_{l},\mathbf{1}_{l}$) $\triangleright$ $v^{0}$: public tree root;  $\mathbf{1}_{l}$: a $l$-dimensional all-ones vector
7:end for
8:function Traverse($v,\mathbf{h}_{1},\mathbf{h}_{2},\mathbf{b},\mathbf{p}_{1},\mathbf{p}_{2}$)
9:  if $v$ is terminal public node then
10:   $\mathbf{u}_{1}$, $\mathbf{u}_{2}$] $\leftarrow\textsc{GetUtility}(v,\mathbf{h}_{1},\mathbf{h}_{2},\mathbf{b})$ $\triangleright$ GetUtility: Compute utilities from a public node and card tensors
11:   return $\mathbf{p}_{2}\circ\mathbf{u}_{1}$, $\mathbf{p}_{1}\circ\mathbf{u}_{2}$] $\triangleright\circ$” denotes element-wise multiplication
12:  end if
13:  $\mathbf{r}_{1}\leftarrow\mathbf{0}_{l}$;  $\mathbf{r}_{2}\leftarrow\mathbf{0}_{l}$;  $\mathbf{R}\leftarrow\mathbf{0}_{m}$ $\triangleright$ zero vectors of length $l$ or $m$
14:  for all $i\in\{1,2\}$ do
15:   if $v$ is player $i$’s public node then
16:     $\mathbf{x}_{i}\leftarrow$ HandTensor($v,\mathbf{h}_{i},\mathbf{b}$) $\triangleright$ HandTensor: Truncate card tensors to $v$’s betting round; construct $l$-batch hand tensors (see Appendix D.1)
17:     $\mathbf{\Phi}_{\cdot,\{j_{1},\dots,j_{l}\}}$, _\_] $\leftarrow\textsc{HandEbdNet}(\mathbf{x}_{i})$
18:     for all $a\in A(J(v))$ do
19:      $\sigma_{original}\leftarrow(\mathbf{\Phi}_{\cdot,\{j_{1},\dots,j_{l}\}})^{\top}\cdot\sigma_{advisor}(v)[a]$
20:      $\mathbf{p}_{i}^{\prime}=\mathbf{p}_{i}\circ\sigma_{original}$;  $\mathbf{p}_{3-i}^{\prime}=\mathbf{p}_{3-i}$
21:      $\mathbf{r}_{1}^{\prime}$, $\mathbf{r}_{2}^{\prime}$] $\leftarrow\textsc{Traverse}(\textsc{NextNode}(v,a),\mathbf{h}_{1},\mathbf{h}_{2},\mathbf{b},\mathbf{p}_{1}^{\prime},\mathbf{p}_{2}^{\prime})$ $\triangleright$ NextNode: Next public node after executing action $a$ at $v$
22:      $\mathbf{r}_{i}\leftarrow\mathbf{r}_{i}+\mathbf{r}_{i}^{\prime}\circ\sigma_{original}$;  $\mathbf{r}_{3-i}\leftarrow\mathbf{r}_{3-i}+\mathbf{r}_{3-i}^{\prime}$
23:      $\tilde{R}_{advisor}(v)[a]\leftarrow\frac{T-1}{T}\tilde{R}_{advisor}(v)[a]+\frac{1}{T}\mathbf{\Phi}_{\cdot,\{j_{1},\dots,j_{l}\}}\cdot(\mathbf{r}_{i}^{\prime}\circ\sigma_{original})$
24:      $\mathbf{R}\leftarrow\mathbf{R}+\textsc{TensorClamp}(\tilde{R}_{advisor}(v)[a],\varepsilon,\infty)$ $\triangleright$ TensorClamp: values $\varepsilon\to\varepsilon$, otherwise unchanged
25:     end for
26:     for all $a\in A(J(v))$ do
[…]
28:      $\bar{\sigma}_{advisor}(v)[a]\leftarrow\frac{T-1}{T}\bar{\sigma}_{advisor}(v)[a]+\frac{1}{T}\sigma_{advisor}(v)[a]$
29:     end for
30:   end if
31:  end for
32:  return $\mathbf{r}_{1}$, $\mathbf{r}_{2}$]
33:end function
34:function HandEbdNet($\mathbf{x}_{q}]_{q\in Q}$) $\triangleright$ $\mathbf{x}_{q}]_{q\in Q}$: $|Q|$-batch hand tensor; $\mathbf{x}_{q}$ as described in Section 4.2
35:  return $\mathbf{\Phi}_{\cdot,q}]_{q\in Q}$, $\mathbf{y}_{q}]_{q\in Q}$] $\triangleright$ $\mathbf{\Phi}_{\cdot,q}$ and $\mathbf{y}_{q}$ as described in Section 4.2
36:end function
Appendix C Theoretical Proofs
Lemma 1 ((lanctot2009monte), Lemma 7).
@@ 2. Large $S^{T}$: Uncontrolled Polynomial Decay (Fractional Scaling)
Appendix D Detailed Implementation of HandEbdNet in Numeral211 Hold’em
D.1 The Construction of the Hand Tensor $\mathbf{x}_{q}$
@@
D.2 The Construction of the Hand Strength Tensor $\mathbf{y}_{q}$
Let’s formalize the construction process of the hand strength tensor $\mathbf{y}$ of a hand $x$. We largely draw on the ideas by fu2024signal to record the strengths of $x$ and its predecessors in their respective rounds.